OpenStudy (anonymous):

Find the area of a circle circumscribed about an equilateral triangle whose side is 18 inches long. i think the answer is 108 pi correct me if im wrong?

OpenStudy (zale101):

you are correct!

OpenStudy (zale101):

to find the Circle's area, you need to use this formula \[\pi r^2 = \pi L^2/3\]

OpenStudy (zale101):

if you want to solve the radius, use the 30-60-90 rule. I'm sure you already know what the radius should look like.